You are a lab technician and need to put together 13 ounces of a 10.64% salt solution. You have stock solution in quart bottles of 6% salt solution and 16% salt solution. Your job is to calculate the number of ounces, accurate to the nearest hunderdth, of each stock solution that is needed to make the required amount.
Answer by checkley77(12844)

.16x+.06(13-x)=.1064*13
.16x+.78-.06x=1.3832
.10x=1.3832-.78
.13x=.6032
x=.6032/.10
x=6.032 oz. of the 16% solution is used.
13-6.032=6.968 oz. of the 6% solution is used.
Proof:
.16*6,032+.06*6.968=.1064*13
.96512+.41808=1.3832
1.3832=1.3832